# Eliciting Latent Predictions from Transformers with the Tuned Lens

### v5

@@ 5.1 Extending Overthinking the Truth
We start by extending a recent use case of the logit lens. Halawi et al. (2023) apply the logit lens to downstream tasks with few-shot prompts, and find that predictions extracted from earlier layers are often more robust to incorrect demonstrations than final layer predictions. Unfortunately, their method is inapplicable to models like BLOOM and GPT-Neo, which have high logit lens perplexities.
We replicate Halawi et al. (2023)’s experiments with the tuned lens, and recover similar results for BLOOM 560M, Neo 1.3B, and Neo 2.7B (Figure 10). Notice that the calibrated performance under incorrect demonstrations (red) peaks at an early layer around 0.4-0.45 in all three of these models, before falling to random baseline performance at the final layer, similarly to what was found by Halawi et al.
@@ 5.2 Detecting Prompt Injections
Given the results from Halawi et al. (2023) and in Figure 10, we hypothesize that the prediction trajectory of the tuned lens on anomalous inputs should be different from the trajectories on normal inputs, and that this could be used to detect anomalous inputs.
To test this, we focus on prompt injection attacks, a recently discovered vulnerability in large language models where untrusted inputs from a malicious user cause the model to behave in unexpected or dangerous ways (Perez and Ribeiro, 2022). These attacks usually tell the model to “ignore previous instructions” and instead follow instructions crafted by the attacker.
To simulate prompt injections, we adapted the attack from Perez and Ribeiro (2022) and ran Pythia 12B on nine multiple choice tasks, both with and without the following adversarial message appended to the few-shot prompt:
@@
We record the tuned prediction trajectory for each data point– that is, for each layer, we record the log probability assigned by the model to each possible answer. [^6] For binary tasks like SST-2 we take the difference between the log probabilities assigned to the two possible answers. We then flatten these trajectories into feature vectors and feed them into two standard outlier detection algorithms: isolation forest (iForest) (Liu et al., 2008) and local outlier factor (LOF) (Breunig et al., 2000), both implemented in scikit-learn (Pedregosa et al., 2011) with default hyperparameters.
Baseline. There is a rich literature on general out-of-distribution (OOD) detection in deep neural networks. One simple technique is to fit a multivariate Gaussian to the model’s final layer hidden states on the training set, and flag inputs as OOD if a new hidden state is unusually far from the training distribution as measured by the Mahalanobis distance (Lee et al., 2018; Mahalanobis, 1936).
Recently, Bai et al. (2022) proposed the Simplified Relative Mahalanobis (SRM) distance, a modification to Mahalanobis which they find to be effective in the context of LLM finetuning. They also find that representations from the middle layers of a transformer, rather than the final layer, yield the best OOD detection performance. We use the SRM at the middle layer as a baseline in our experiments.
Experimental setup. We fit each anomaly detection model exclusively on prediction trajectories from normal prompts without prompt injections, and evaluate them on a held out test set containing both normal and prompt-injected trajectories. This ensures that our models cannot overfit to the prompt injection distribution. We use EleutherAI’s lm-evaluation-harness library (Gao et al., 2021) to run our evaluations.
@@ 5.3 Measuring Example Difficulty
Early exiting strategies like CALM (Schuster et al., 2022) and DeeBERT (Xin et al., 2020) are based on the observation that “easy” examples require less computation to classify than “difficult” examples. If an example is easy, the model should quickly converge to the right answer in early layers, making it possible to skip the later layers without a significant drop in prediction quality. Conversely, the number of layers needed to converge on an answer can be used to measure the difficulty of an example.
We propose to use the tuned lens to estimate example difficulty in pretrained transformers, without the need to finetune the model for early exiting. Following Baldock et al. (2021)’s work on computer vision models, we define the prediction depth of a prompt $\boldsymbol{x}$ to be the number of layers after which a model’s top-1 prediction for $\boldsymbol{x}$ stops changing.
To validate the prediction depth, we measure its correlation with an established difficulty metric: the iteration learned. The iteration learned is defined as the earliest training step $\tau$ where the model’s top-1 prediction for a datapoint $\boldsymbol{x}$ is fixed (Toneva et al., 2018). Intuitively, we might expect that examples which take a long time to learn during training would tend to require many layers of computation to classify at inference time. Baldock et al. (2021) indeed show such a correlation, using k-NN classifiers to elicit early predictions from the intermediate feature maps of image classifiers.
@@ 6 Discussion
Figure 11: Token-level prediction depths for Pythia 12B computed on the abstract of OpenAI (2023). Warm colors have high prediction depth, while cool colors indicate low depth.
@@ Appendix B Qualitative Results
Figure 16: Tuned lens prediction trajectory for Pythia 12B prompted with the abstract of Brown et al. [2020].
@@ Appendix C Transformers Perform Iterative Inference
Jastrzębski et al. [2017] argue that skip connections encourage neural networks to perform “iterative inference” in the following sense: each layer reliably updates the hidden state in a direction of decreasing loss. While their analysis focused on ResNet image classifiers, their theoretical argument applies equally well to transformers, or any neural network featuring residual connections. We reproduce their theoretical analysis below.
@@
To empirically measure the extent to which residuals do align with the negative gradient, Jastrzębski et al. [2017] compute the cosine similarity between $F(\boldsymbol{h}_{i})$ and $\frac{\partial\mathcal{L}(\boldsymbol{h}_{i})}{\partial\boldsymbol{h}_{i}}$ at each layer of a ResNet image classifier. They find that it is consistently negative, especially in the final stage of the network.
@@ C.1 Zero-shot robustness to layer deletion
Stochastic depth [Huang et al., 2016], also known as LayerDrop [Fan et al., 2019], is a regularization technique which randomly drops layers during training, and can be applied to both ResNets and transformers. Veit et al. [2016] show that ResNets are robust to the deletion of layers even when trained without stochastic depth, while CNN architectures without skip connections are not robust in this way. These results strongly support the idea that adjacent layers in a ResNet encode fundamentally similar representations [Greff et al., 2016].
To the best of our knowledge, this experiment had never been replicated before in transformers. We did so and report our results in Figure 19 above. We find that only the very first layer is crucial for performance; every other layer ablation induces a nearly imperceptible increase in perplexity. Interestingly, Veit et al. [2016] also found that the first layer is exceptionally important in ResNets, suggesting that this is a general phenomenon.
@@ D.1 Ablation techniques
This “zeroes out” a part of $\boldsymbol{x}$, ensuring that $\langle\boldsymbol{x}^{\prime},\boldsymbol{u}\rangle=\boldsymbol{0}$ for all $\boldsymbol{u}$ in $B$. But recent work has pointed out that setting activations to zero may inadvertently push a network out of distribution, making the experimental results hard to interpret [Wang et al., 2022, Chan et al., 2022].
@@ Appendix E Static interpretability analysis
Several interpretability methods aim to analyze parameters that “write" (in the sense of Elhage et al. [2021]) to intermediate hidden states of a language model [Millidge and Black, 2022, Dar et al., 2022, Geva et al., 2022, 2020], and even use this analysis to successfully edit model behavior [Geva et al., 2022, Millidge and Black, 2022, Dai et al., 2022]. Given that the tuned lens aims to provide a “less biased" view of intermediate hidden states, we should expect the tuned lens to preserve their effectiveness. We test both static analysis and model editing methods, and confirm this hypothesis. With static analysis, the tuned lens appears to never decrease performance, and for some models, even increases their performance. With model editing, we found the tuned lens to outperform the logit lens on OPT-125m and perform equivalently on other models for the task of toxicity reduction.
@@ E.1 Static Analysis
Many parameters in transformer language models have at least one dimension equal to that of their hidden state, allowing us to multiply by the unembedding matrix to project them into token space. Surprisingly, the resulting top $k$ tokens are often meaningful, representing interpretable semantic and syntactic clusters, such as “prepositions” or “countries” [Millidge and Black, 2022, Dar et al., 2022, Geva et al., 2022]. This has successfully been applied to the columns of the MLP output matrices [Geva et al., 2022], the singular vectors of the MLP input and output matrices [Millidge and Black, 2022], and the singular vectors of the attention output-value matrices $W_{OV}$ [Millidge and Black, 2022].
In short [^7] The exact interpretation depends on the method used - for instance, the rows of $W_{\text{out}}$ can be interpreted as the values in a key-value memory store [Geva et al., 2020]., we can explain this as: many model parameters directly modify the model’s hidden state [Elhage et al., 2021], and when viewed from the right angle, the modifications they make are often interpretable. These “interpretable vectors” occur significantly more often than would be expected by random chance [Geva et al., 2022], and can be used to edit the model’s behaviour, like reducing probability of specific tokens [Millidge and Black, 2022], decreasing toxicity [Geva et al., 2022], or editing factual knowledge [Dai et al., 2022].
@@
To test this hypothesis, we developed a novel automated metric for evaluating the performance of these parameter interpretability methods, based on the pretrained BPEmb encoder [Heinzerling and Strube, 2018], enabling much faster and more objective experimentation than possible with human evaluation. We describe this method in Appendix E.3.
@@
Table 3: The mean interpretability scores (as measured in Section˜E.3) for Pythia 125M, with several different interpretability techniques [Millidge and Black, 2022, Geva et al., 2020], comparing both the tuned lens and logit lens to randomly generated matrices. Where applicable, the notation (L) and (R) indicates that the results are for the left and right singular vectors, respectively.
@@ E.2 Interpretability score distributions
Figure 21: The interpretability scores for the right singular vectors of the OV matrices $W_{OV}$ (following Millidge and Black [2022]) in Pythia 125M, compared with a randomly generated matrix, for both the logit lens (left), and the tuned lens (right).
@@ E.3 Automated interpretability analysis method
While humans can easily tell if a list of words represents some coherent category, this can be subjective and time-consuming to evaluate manually. Some previous work has attempted to automate this process, using GPT-3 in Millidge and Black [2022] or the Perspective API in Geva et al. [2022], but these can still be too slow for quick experimentation.
We created our own method, motivated by the observation that much of the human interpretable structure of these tokens appears to be related to their mutual similarity. Then, the mutual similarity of words can easily be measured by their cosine similarity under a word embedding [Sitikhu et al., 2019]. In particular, we use the pretrained BPEmb encoder [Heinzerling and Strube, 2018]  [^9] We also tried the fastText encoder [Bojanowski et al., 2016], but found BPEmb to give results better matching human judgment.. We encode each token, measure the cosine similarity pairwise between all tokens, and take the average value:
@@ E.4 Improved Model Editing
Geva et al. [2022] edit GPT-2-medium to be less toxic by changing the coefficients which feed into the columns of the MLP output matrix, under the key-value framework developed in Geva et al. [2020]. They define both the value vector $v_{i}^{\ell}$ (i.e. the columns of the MLP’s $W_{\text{out}}$) and a coefficient $m_{i}^{\ell}$ (i.e. the output of $W_{\text{in}}$ before the activation function) for each column $i$ and each layer $\ell$. This is used to reduce the toxicity of GPT-2-medium by editing the coefficients of specific value vectors. They find:
@@
They then concatenate these top-30 tokens for every layer and column, sending it to Perspective API, a Google API service which can return a toxicity score. They then sampled from the vectors that scored < 0.1 toxicity (where a score > 0.5 is classified as toxic), set each value vector’s respective coefficient to 3 (because 3 was a higher than average activation), and ran the newly edited model through a subset of REALTOXICPROMPTS [Gehman et al., 2020] to measure the change in toxicity compared to the original model.
@@ E.5 Implementation
The initial implementation of many of our static analysis experiments was helped by the use of the transformer_lens library [Nanda, 2022].

### v6

@@ 5.2 Extending Overthinking the Truth
halawi2023overthinking apply the logit lens to downstream tasks with few-shot prompts, and find that predictions extracted from earlier layers are often more robust to incorrect demonstrations than final layer predictions. Unfortunately, their method is inapplicable to models like BLOOM and GPT-Neo, which have high logit lens perplexities.
We replicate halawi2023overthinking’s experiments with the tuned lens, and recover similar results for BLOOM 560M, Neo 1.3B, and Neo 2.7B (Figure 11). Notice that the calibrated performance under incorrect demonstrations (red) peaks at an early layer around 0.4-0.45 in all three of these models, before falling to random baseline performance at the final layer, similarly to what was found by halawi2023overthinking
@@ 5.3 Detecting Prompt Injections
Given the results from halawi2023overthinking and in Figure 11, we hypothesize that the prediction trajectory of the tuned lens on anomalous inputs should be different from the trajectories on normal inputs, and that this could be used to detect anomalous inputs.
To test this, we focus on prompt injection attacks, a recently discovered vulnerability in large language models where untrusted inputs from a malicious user cause the model to behave in unexpected or dangerous ways (perez2022ignore). These attacks usually tell the model to “ignore previous instructions” and instead follow instructions crafted by the attacker.
To simulate prompt injections, we adapted the attack from perez2022ignore and ran Pythia 12B on nine multiple choice tasks, both with and without the following adversarial message appended to the few-shot prompt:
@@
We record the tuned prediction trajectory for each data point– that is, for each layer, we record the log probability assigned by the model to each possible answer. [^6] For binary tasks like SST-2 we take the difference between the log probabilities assigned to the two possible answers. We then flatten these trajectories into feature vectors and feed them into two standard outlier detection algorithms: isolation forest (iForest) (liu2008isolation) and local outlier factor (LOF) (breunig2000lof), both implemented in scikit-learn (scikit-learn) with default hyperparameters.
Baseline. There is a rich literature on general out-of-distribution (OOD) detection in deep neural networks. One simple technique is to fit a multivariate Gaussian to the model’s final layer hidden states on the training set, and flag inputs as OOD if a new hidden state is unusually far from the training distribution as measured by the Mahalanobis distance (lee2018simple; mahalanobis1936generalized).
Recently, bai2022training proposed the Simplified Relative Mahalanobis (SRM) distance, a modification to Mahalanobis which they find to be effective in the context of LLM finetuning. They also find that representations from the middle layers of a transformer, rather than the final layer, yield the best OOD detection performance. We use the SRM at the middle layer as a baseline in our experiments.
Experimental setup. We fit each anomaly detection model exclusively on prediction trajectories from normal prompts without prompt injections, and evaluate them on a held out test set containing both normal and prompt-injected trajectories. This ensures that our models cannot overfit to the prompt injection distribution. We use EleutherAI’s lm-evaluation-harness library (eval-harness) to run our evaluations.
@@ 5.4 Measuring Example Difficulty
Early exiting strategies like CALM (schuster2022confident) and DeeBERT (xin2020deebert) are based on the observation that “easy” examples require less computation to classify than “difficult” examples. If an example is easy, the model should quickly converge to the right answer in early layers, making it possible to skip the later layers without a significant drop in prediction quality. Conversely, the number of layers needed to converge on an answer can be used to measure the difficulty of an example.
We propose to use the tuned lens to estimate example difficulty in pretrained transformers, without the need to finetune the model for early exiting. Following baldock2021deep’s work on computer vision models, we define the prediction depth of a prompt $\boldsymbol{x}$ to be the number of layers after which a model’s top-1 prediction for $\boldsymbol{x}$ stops changing.
To validate the prediction depth, we measure its correlation with an established difficulty metric: the iteration learned. The iteration learned is defined as the earliest training step $\tau$ where the model’s top-1 prediction for a datapoint $\boldsymbol{x}$ is fixed (toneva2018empirical). Intuitively, we might expect that examples which take a long time to learn during training would tend to require many layers of computation to classify at inference time. baldock2021deep indeed show such a correlation, using k-NN classifiers to elicit early predictions from the intermediate feature maps of image classifiers.
@@ 6 Discussion
Figure 12: Token-level prediction depths for Pythia 12B computed on the abstract of openai2023gpt4. Warm colors have high prediction depth, while cool colors indicate low depth.
@@ Appendix B Qualitative Results
Figure 16: Tuned lens prediction trajectory for Pythia 12B prompted with the abstract of brown2020language.
@@ Appendix C Transformers Perform Iterative Inference
jastrzkebski2017residual argue that skip connections encourage neural networks to perform “iterative inference” in the following sense: each layer reliably updates the hidden state in a direction of decreasing loss. While their analysis focused on ResNet image classifiers, their theoretical argument applies equally well to transformers, or any neural network featuring residual connections. We reproduce their theoretical analysis below.
@@
To empirically measure the extent to which residuals do align with the negative gradient, jastrzkebski2017residual compute the cosine similarity between $F(\boldsymbol{h}_{i})$ and $\frac{\partial\mathcal{L}(\boldsymbol{h}_{i})}{\partial\boldsymbol{h}_{i}}$ at each layer of a ResNet image classifier. They find that it is consistently negative, especially in the final stage of the network.
@@ C.1 Zero-shot robustness to layer deletion
Stochastic depth [huang2016deep], also known as LayerDrop [fan2019reducing], is a regularization technique which randomly drops layers during training, and can be applied to both ResNets and transformers. veit2016residual show that ResNets are robust to the deletion of layers even when trained without stochastic depth, while CNN architectures without skip connections are not robust in this way. These results strongly support the idea that adjacent layers in a ResNet encode fundamentally similar representations [greff2016highway].
To the best of our knowledge, this experiment had never been replicated before in transformers. We did so and report our results in Figure 19 above. We find that only the very first layer is crucial for performance; every other layer ablation induces a nearly imperceptible increase in perplexity. Interestingly, veit2016residual also found that the first layer is exceptionally important in ResNets, suggesting that this is a general phenomenon.
@@ D.1 Ablation techniques
This “zeroes out” a part of $\boldsymbol{x}$, ensuring that $\langle\boldsymbol{x}^{\prime},\boldsymbol{u}\rangle=\boldsymbol{0}$ for all $\boldsymbol{u}$ in $B$. But recent work has pointed out that setting activations to zero may inadvertently push a network out of distribution, making the experimental results hard to interpret [wang2022interpretability, chan2022causal].
@@ Appendix E Static interpretability analysis
Several interpretability methods aim to analyze parameters that “write" (in the sense of elhage2021mathematical) to intermediate hidden states of a language model [millidge2022svd, dar2022analyzing, geva2022transformer, geva2020transformer], and even use this analysis to successfully edit model behavior [geva2022transformer, millidge2022svd, dai-etal-2022-knowledge]. Given that the tuned lens aims to provide a “less biased" view of intermediate hidden states, we should expect the tuned lens to preserve their effectiveness. We test both static analysis and model editing methods, and confirm this hypothesis. With static analysis, the tuned lens appears to never decrease performance, and for some models, even increases their performance. With model editing, we found the tuned lens to outperform the logit lens on OPT-125m and perform equivalently on other models for the task of toxicity reduction.
@@ E.1 Static Analysis
Many parameters in transformer language models have at least one dimension equal to that of their hidden state, allowing us to multiply by the unembedding matrix to project them into token space. Surprisingly, the resulting top $k$ tokens are often meaningful, representing interpretable semantic and syntactic clusters, such as “prepositions” or “countries” [millidge2022svd, dar2022analyzing, geva2022transformer]. This has successfully been applied to the columns of the MLP output matrices [geva2022transformer], the singular vectors of the MLP input and output matrices [millidge2022svd], and the singular vectors of the attention output-value matrices $W_{OV}$ [millidge2022svd].
In short [^7] The exact interpretation depends on the method used - for instance, the rows of $W_{\text{out}}$ can be interpreted as the values in a key-value memory store [geva2020transformer]., we can explain this as: many model parameters directly modify the model’s hidden state [elhage2021mathematical], and when viewed from the right angle, the modifications they make are often interpretable. These “interpretable vectors” occur significantly more often than would be expected by random chance [geva2022transformer], and can be used to edit the model’s behaviour, like reducing probability of specific tokens [millidge2022svd], decreasing toxicity [geva2022transformer], or editing factual knowledge [dai-etal-2022-knowledge].
@@
To test this hypothesis, we developed a novel automated metric for evaluating the performance of these parameter interpretability methods, based on the pretrained BPEmb encoder [heinzerling2018bpemb], enabling much faster and more objective experimentation than possible with human evaluation. We describe this method in Appendix E.3.
@@
Table 3: The mean interpretability scores (as measured in Section˜E.3) for Pythia 125M, with several different interpretability techniques [millidge2022svd, geva2020transformer], comparing both the tuned lens and logit lens to randomly generated matrices. Where applicable, the notation (L) and (R) indicates that the results are for the left and right singular vectors, respectively.
@@ E.2 Interpretability score distributions
Figure 21: The interpretability scores for the right singular vectors of the OV matrices $W_{OV}$ (following millidge2022svd) in Pythia 125M, compared with a randomly generated matrix, for both the logit lens (left), and the tuned lens (right).
@@ E.3 Automated interpretability analysis method
While humans can easily tell if a list of words represents some coherent category, this can be subjective and time-consuming to evaluate manually. Some previous work has attempted to automate this process, using GPT-3 in millidge2022svd or the Perspective API in geva2022transformer, but these can still be too slow for quick experimentation.
We created our own method, motivated by the observation that much of the human interpretable structure of these tokens appears to be related to their mutual similarity. Then, the mutual similarity of words can easily be measured by their cosine similarity under a word embedding [Sitikhu2019ACO]. In particular, we use the pretrained BPEmb encoder [heinzerling2018bpemb]  [^9] We also tried the fastText encoder [bojanowski2016enriching], but found BPEmb to give results better matching human judgment.. We encode each token, measure the cosine similarity pairwise between all tokens, and take the average value:
@@ E.4 Improved Model Editing
geva2022transformer edit GPT-2-medium to be less toxic by changing the coefficients which feed into the columns of the MLP output matrix, under the key-value framework developed in geva2020transformer. They define both the value vector $v_{i}^{\ell}$ (i.e. the columns of the MLP’s $W_{\text{out}}$) and a coefficient $m_{i}^{\ell}$ (i.e. the output of $W_{\text{in}}$ before the activation function) for each column $i$ and each layer $\ell$. This is used to reduce the toxicity of GPT-2-medium by editing the coefficients of specific value vectors. They find:
@@
They then concatenate these top-30 tokens for every layer and column, sending it to Perspective API, a Google API service which can return a toxicity score. They then sampled from the vectors that scored < 0.1 toxicity (where a score > 0.5 is classified as toxic), set each value vector’s respective coefficient to 3 (because 3 was a higher than average activation), and ran the newly edited model through a subset of REALTOXICPROMPTS [gehman-etal-2020-realtoxicityprompts] to measure the change in toxicity compared to the original model.
@@ E.5 Implementation
The initial implementation of many of our static analysis experiments was helped by the use of the transformer_lens library [nandatransformerlens2022].